OpenStudy (anonymous):

true or false: log 1/2 = log 1 / log 2

OpenStudy (anonymous):

False. log(1/2)=log(1)-log(2)

OpenStudy (anonymous):

true or false log 4x^3 = log 4+3 log x

OpenStudy (zarkon):

\[\log(4x^3)=\log(4)+\log(x^3)=\log(4)+3\log(x)\]